ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 236, 2020

Aka, M., M. Einsiedler, H. Li and A. Mohammadi On effective equidistribution for quotients of SL(d, ℝ)

365

Alpay, D., I. L. Paiva and D. C. Struppa A general setting for functions of Fueter variables: differentiability, rational functions, Fock module and related topics

207

Anoop, V. P. and S. Parui The Hardy inequality and fractional Hardy inequality for the Dunkl Laplacian

247

Araújo, G., P. H. Enflo, G. A. Muñoz-Fernández, D. L. Rodríguez-Vidanes and J. B. Seoane-Sepúlveda Quantitative and qualitative estimates on the norm of products of polynomials

727

Bárcenas, N. and J. Cantarero A completion theorem for fusion systems

501

Belliart, M. A classification of certain codimension one locally free actions of nilpotent Lie groups up to a differentiable orbital conjugacy

279

Black, A., U. Cetin, F. Frick, A. Pacun and L. Setiabrata Fair splittings by independent sets in sparse graphs

603

Boney, W. Model theoretic characterizations of large cardinals  133
Campos, M. On the number of sets with a given doubling constant  711

Cantarero, J. See N. Bárcenas Cetin, U. See A. Black Chernousov, V. I., A. S. Rapinchuk and I. A. Rapinchuk The

finiteness of the genus of a finite-dimensional division algebra, and some generalizations

747

Chlebicka, I. Regularizing effect of the lower-order terms in elliptic problems with Orlicz growth

967

Chyzhykov, I. E. Asymptotic behaviour of pth means of analytic and subharmonic functions in the unit disc and angular distribution of zeros

931

Cubides Kovacsics, P. and F. Delon Definable functions in tame expansions of algebraically closed valued fields

651

Delon, F. See P. Cubides Kovacsics Dobrowolski, J. and F. O. Wagner On ω-categorical groups and

rings of finite burden 801 Einsiedler, M. See M. Aka Enflo, P. H. See G. Araújo Frick, F. See A. Black Goodrich, C. and C. Lizama A transference principle for nonlocal

operators using a convolutional approach: fractional monotonicity and convexity

533

Grupel, U. Sampling by intersections with random geodesics  423

Hachtman, S. and D. Sinapova The super tree property at the successor of a singular

473

Hagen, M. F. and T. Susse On hierarchical hyperbolicity of cubical groups

45

Hayut, Y. and A. Karagila Critical cardinals  449

Jiang, Z. and A. Polyanskii Forbidden subgraphs for graphs of bounded spectral radius, with applications to equiangular lines

393

Kahn, J. and J. Park The number of 4-colorings of the Hamming cube

629

Karagila, A. See Y. Hayut Kitture, R. D. See T. K. Naik

Lei, A. and F. Sprung Ranks of elliptic curves over ℤ_{p}^{2}-extensions  183

Letzter, S. and J. Sahasrabudhe On existentially complete triangle-free graphs

591

Li, H. See M. Aka Lizama, C. See C. Goodrich

Lücke, P. Partition properties for simply definable colourings  841

Mathieu, M. and P. Tradacete Strictly singular multiplication operators on L(X)

685

Melleray, J. Dynamical simplices and Borel complexity of orbit equivalence

317

Mohammadi, A. See M. Aka Muñoz-Fernández, G. A. See G. Araújo Naik, T. K., R. D. Kitture and M. K. Yadav Finite p-groups of

nilpotency class 3 with two conjugacy class sizes

899

Pacun, A. See A. Black Paiva, I. L. See D. Alpay Pantone, J. and V. Vatter Growth rates of permutation classes:

Categorization up to the uncountability threshold

1

Paolini, G. and S. Shelah On a cardinal invariant related to the Haar measure problem

305

Park, J. See J. Kahn Parui, S. See V. P. Anoop Petrov, F. and C. Pohoata Improved bounds for progression-free

sets in C_{8}^{n} 345

Pohoata, C. See F. Petrov Polyanskii, A. See Z. Jiang

Puglisi, D. A property for the Monge–Ampère equation  959

Rapinchuk, A. S. See V. I. Chernousov Rapinchuk, I. A. See V. I. Chernousov Rodríguez-Vidanes, D. L. See G. Araújo Sahasrabudhe, J. See S. Letzter Seoane-Sepúlveda, J. B. See G. Araújo Setiabrata, L. See A. Black Shelah, S. See G. Paolini Sinapova, D. See S. Hachtman Sprung, F. See A. Lei Struppa, D. C. See D. Alpay Susse, T. See M. F. Hagen Tanaka, H. Perturbed finite-state Markov systems with holes and

Perron complements of Ruelle operators

91

Tradacete, P. See M. Mathieu Vatter, V. See J. Pantone Wagner, F. O. See J. Dobrowolski Yadav, M. K. See T. K. Naik